# Convergence Analysis of Block Newton Methods

### v1

@@ 1 Introduction
where $\sigma(t)=\max\{0,t\}$ is the ReLU activation function. Denote by ${\bf c}=(c_{0},\dots,c_{n})^{T}\in\mathbb{R}^{n+1}$ the linear parameters and by ${\bf b}=(b_{1},\dots,b_{n})^{T}\in\mathbb{R}^{n}$ the nonlinear parameters, and denote all parameters by $\mbox{${\theta}$}=(c_{0},c_{1},\dots,c_{n},b_{1},\dots,b_{n})^{T}\in\mathbb{R}^{2n+1}$. Notice that we chose all hidden layer weights to be one by normalization. Each function in ${\cal M}_{n}(I)$
@@
depends on the parameters $\bf c$} and $\bf b$} and is a piecewise linear function of $x\in[0,1]$ with the breaking (mesh) points $\left\{b_{i}\right\}_{i=1}^{n}$.
It is well-known that the set ${\cal M}_{n}(I)$ is equivalent to the free knot splines (FKS) (see [Schumaker]), where FKS uses local hat basis functions. FKS can substantially enhance the approximation order and reduce the number of degrees of freedom for non-smooth functions (see [Burchard74] and the discussion in [cai2024fast]). However, determining the optimal knot locations (the nonlinear parameters $\bf b$}) results in a high-dimensional, non-convex optimization problem that is computationally intensive.
Recently, for one-dimensional elliptic differential equations, a damped block Newton (dBN) method was developed in [cai2024fast, cai2024fast2] for solving the shallow Ritz discretization in order to efficiently move the mesh points. The dBN method adopts a classical outer-inner iterative method (see, e.g., [Ainsworth2020PlateauPI, Ainsworth2022, pmlr-v107-cyr20a, Golub1973, separable3, park2022neuron, separable2]), alternating between updates of the linear and nonlinear parameters. While this approach is natural, it does not resolve the key challenges. First, optimizing the linear parameters requires solving dense, ill-conditioned linear systems. Second, optimizing the nonlinear parameters leads to a system of nonlinear equations, where the Hessian matrix may be singular. Effective strategies for addressing these challenges were presented in [cai2024fast, cai2024fast2], resulting in an efficient iterative method for finding a nearly optimal neural network approximation for one-dimensional elliptic differential equations and least-squares function approximation problems. For multi-dimensional least-squares problems using shallow ReLU NN, a structure-guided Gauss-Newton method was studied in [Cai24GN], where the singularities of the resulting Gauss-Newton matrix was removed to design a much faster solver than the popular Levenberg-Marquardt method [Levenberg, Marquardt] using heuristic regularization or shifting techniques.
@@
The purpose of this paper is to analyze local convergence of block Newton (BN) methods for one-dimensional shallow neural network approximation to functions and diffusion-reaction problems. The BN method consists of the $2\times 2$ block nonlinear Gauss-Seidel, linear Gauss-Seidel, or Jacobi method for outer iteration and the Newton method for inner iteration. The blocks are corresponding to the linear and the nonlinear parameters. By following the machinery in [ortega1970iterative] on local convergence of the componentwise Gauss-Seidel method, we first develop a local convergence theory for the block Newton methods applicable to shallow ReLU NNs in both one and multiple dimensions, provided that the Hessian matrix at a critical point is symmetric positive definite (SPD) and that the $2\times 2$ block nonlinear Gauss-Seidel, linear Gauss-Seidel, or Jacobi matrix is invertible. For one-dimensional diffusion–reaction problems and least-squares function approximation, we then derive sufficient conditions on SPD of the Hessian. Unlike the common optimization methods, the reduced BN (rBN) method allows reduction of the number of parameters during the optimization process when some neurons have no contribution in approximation or are at nearly optimal locations.

### v2

@@ 1 Introduction
where $\sigma(t)=\max\{0,t\}$ is the ReLU activation function. Denote by ${\bf c}=(c_{0},\dots,c_{n})^{T}\in\mathbb{R}^{n+1}$ and ${\bf b}=(b_{1},\dots,b_{n})^{T}\in\mathbb{R}^{n}$ the respective linear parameters and nonlinear parameters, and denote by $\mbox{${\theta}$}=(c_{0},c_{1},\dots,c_{n},b_{1},\dots,b_{n})^{T}\in\mathbb{R}^{2n+1}$ all parameters. Notice that the weights of all hidden layer is chosen to be one by normalization (see [Liu_Cai_Chen_2022]). Each function in ${\cal M}_{n}(I)$
@@
depends on the parameters $\bf c$} and $\bf b$} and is a piecewise linear function of $x\in[0,1]$ with respect to the partition by the breaking (mesh) points: $0=b_{0}<b_{1}<\cdots<b_{n}<b_{n+1}=1$.
It is known (see, e.g., [Liu_Cai_Chen_2022]) that the set ${\cal M}_{n}(I)$ is equivalent to the free-knot splines (FKS) (see [Schumaker]), where FKS utilizes local hat basis functions. FKS can substantially enhance the approximation order and reduce the number of degrees of freedom for non-smooth functions (see [Burchard74] and the discussion in [cai2024fast]). As an example of this, the order of the best approximation to $f(x)=x^{\alpha}$ ($0<\alpha<1$) on $I$ is merely $\alpha<1$ (i.e., ${\cal O}(n^{-\alpha})$) when using finite elements on a fixed uniform mesh, whereas for FKS, the order becomes one (i.e., ${\cal O}(n^{-1})$) no matter how small the exponent $\alpha>0$ is (see, e.g., [daubechies, Schumaker]). This is a huge improvement in approximation.
Despite the remarkable approximation capability of FKS for non-smooth functions, there are two essential difficulties that have led numerical analysts moving away from FKS: (1) no successful extension of FKS to two or higher dimensions has been achieved, and (2) determining the optimal knot locations (the nonlinear parameters $\mathbf{b}$) results in a high-dimensional, non-convex optimization problem, that is computationally expensive and hence dismisses its benefit in approximation. To the best of our knowledge, there are no available efficient optimization schemes that enable the competitiveness of FKS. While the first issue may be addressed by employing neural networks due to their global supported basis functions, the second issue still remains a major challenge.
To address this challenge, a major advance on fast iterative solver was recently made in [cai2024fast, cai2024fast2] for solving non-convex optimization problems arising from shallow ReLU NN approximation to a given function or solutions of elliptic differential equations in one dimension. Specifically, a well-designed damped block Newton (dBN) method was developed. First, the dBN adopts a classical outer-inner iterative strategy (see, e.g., [Ainsworth2020PlateauPI, Ainsworth2022, pmlr-v107-cyr20a, Golub1973, separable3, park2022neuron, separable2]), alternating between updates of the linear and nonlinear parameters. Second, to solve the resulting dense, ill-conditioned linear systems due to the global basis functions of NNs, the dBN uses the fact that the exact inversion of those matrices can be represented in terms of products of sparse matrices (see [cai2024fast2]). Third, the dBN deals directly with singularities of the Hessian for the nonlinear parameters by removing neurons whose linear parameters are small or whose nonlinear parameters have reached nearly optimal locations. As a result, the computational cost per iteration of the dBN is ${\cal O}(n)$, and numerical experiments show that this method is capable of moving mesh points effectively and efficiently. Beyond strong one dimensional results, the methodology of the dBN is conceptually promising for higher dimensions, as it demonstrates how to design iterative solvers that exploit the problem structure and the approximation and geometric meanings of the NN parameters.
@@
The purpose of this paper is to provide a theoretical guarantee on why this sophisticated dBN moves the mesh points efficiently. This will be done by analyzing local convergence of block Newton (BN) methods. The BN method consists of the $2\times 2$ block nonlinear Gauss-Seidel, linear Gauss-Seidel, or Jacobi method for outer iteration and the Newton method for inner iteration. The blocks are corresponding to the linear and the nonlinear parameters. By following the machinery in [ortega1970iterative] on local convergence of the componentwise Gauss-Seidel method, we first develop a local convergence theory for the block Newton methods applicable to shallow ReLU NNs in both one and multiple dimensions, provided that the Hessian matrix at a critical point is symmetric positive definite (SPD) and that the $2\times 2$ block nonlinear Gauss-Seidel, linear Gauss-Seidel, or Jacobi matrix is invertible.
By expressing the BN method as a fixed-point iteration, local convergence of the BN method is established by showing that a norm of the corresponding Jacobian at the critical point is strictly less than one. Note that derivation of the Jacobian matrix is non-trivial (see Lemma 3.3). To guarantee feasibility of each Newton step, the BN method is modified to the reduced BN (rBN) method that allows a reduction in the number of parameters during the optimization process. Local convergence of the rBN is justified by showing that some nonlinear parameters are at nearly optimal locations.